OpenStudy (anonymous):

log(3x+1)=2 @jim_thompson5910 i'll show you the work

pooja195 (pooja195):

base it, so it is 10^log(3x+1) = 10^2 the 10^log(base 10) cancels because they are inverse functions and thus you are left with: 3x+1=100 3x=99 x=33 If you want to check, punch in log(3*33 +1) log100=2

OpenStudy (anonymous):

Thank you very much that was very helpful =) @pooja195

pooja195 (pooja195):

no problem